Find the distance between each pair of points whose coordinates are given Express answers in simplest radical form and as decimal approximations to the nearest hundreth
1.(9,7),(1,1)
2.(5,2),(8,-2)
3.(1,-3, (1,4)

Answer by kiru_khandelwal(79)

Distance between points with coordinates (x1,y1) and (x2,y2) is sqrt((x2-x1)*(x2-x1) + (y2-y1)*(y2-y1))
Therefore distance between
1) (9,7),(1,1)
=> x2-x1 = 1-9 = -8
=> y2-y1 = 1-7 = -6
Distance between two points = sqrt ((-8)*(-8) + (-6)*(-6))
=> sqrt(64 + 36) => sqrt (100) => 10
Similarly distancex between other pairs can also be found out